OpenStudy (anonymous):

Use the distributive property to find the product of the following: 5(x+5)= (Simplify Answer)

OpenStudy (anonymous):

@perl

OpenStudy (perl):

the distributive property: a( b + c ) = a*b + a*c

OpenStudy (perl):

you can use that as a template

Nnesha (nnesha):

distribute bracket by 5 its mean multiply 5 by every number which is in bracket

OpenStudy (anonymous):

5x+25

OpenStudy (anonymous):

5(x+5)= 25

Nnesha (nnesha):

5(x + 5) is not equal just @educated

Nnesha (nnesha):

just 25***************

OpenStudy (anonymous):

._. Yeah I know, sorry was in a hurry. Should have double checked.